# Algorithmic stability implies training-conditional coverage for distribution-free prediction methods

### v2

@@ 4.3 In-sample, out-of-sample, or neither?
However, there is a third option: the test point might be in the training set for one model, but not the other. In particular, Barber et al. (2021) also consider this alternative type of stability, which (adjusting to the notation of our present paper) can be expressed by the quantity
@@
We write “in/out” in this definition to emphasize the fact that, in this case, the test point $X_{n+1}$ is in the training sample for one of the models ($\widehat{\mu}_{n+1}$) but out of the training sample for the other ($\widehat{\mu}_{n}$). (We remark that Barber et al. (2021) refer to this as “in-sample stability” but we change the terminology here to allow for all the various types of stability under consideration.)
@@ 5.1 Relating $1$-stability and $m$-stability
At a high level, it is intuitive that achieving $m$-stability, for some $m>1$, is more challenging than achieving $1$-stability. For swap-stability, this intuition is correct, as shown in the following proposition (we defer the proof to Section A.5).
@@
For any algorithm $\mathcal{A}$ and distribution $P$, for any $n\geq m^{\prime}>m>0$,
@@
On the other hand, if we define stability by removing/adding data points (rather than by swapping), this may not necessarily hold, due to degenerate counterexamples (e.g., we can take an algorithm $\mathcal{A}$ which, when run on a training sample of size $n$, returns the constant function $\widehat{\mu}(x)\equiv C\cdot\textnormal{mod}(n,2)$ for an arbitrary constant $C>0$; then $\beta_{1,n}^{\textnormal{out}}(\mathcal{A},P)=\beta_{1,n}^{\textnormal{in}}(\mathcal{A},P)=C$, but $\beta_{m,n}^{\textnormal{out}}(\mathcal{A},P)=\beta_{m,n}^{\textnormal{in}}(\mathcal{A},P)=0$ for any even $m$). Nonetheless, in practice, we would generally expect $\beta_{m,n}^{\textnormal{out}}(\mathcal{A},P)$ and $\beta_{m,n}^{\textnormal{in}}(\mathcal{A},P)$ to each be monotone increasing in $m$.
@@
For any $\mathcal{A}$ and any $P$, for any $n,m\geq 0$, the following inequality holds for out-of-sample stability:
@@
Proof of Lemma 5.2.
by definition of out-of-sample $1$-swap-stability (together with the assumption of symmetry of $\mathcal{A}$). Again, a similar argument holds for in-sample swap-stability. ∎
In particular, we will see next that this lemma immediately implies the results of Theorems 3.6 and 3.7. However, as we will demonstrate below both empirically (in Section 5.3) and theoretically (in Section 5.4), this result is fundamentally somewhat weak: it does not capture the favorable stability properties that appear to hold in practice, and we will see that the $m$-stability of algorithms typically tends to be much lower than this coarse upper bound.
5.2 Proofs of asymptotic results
Proof of Theorems 3.6 and 3.7.
Define
$a_{n}=\max_{n^{\prime}\geq n-1}\frac{\beta_{1,n^{\prime}}^{\textnormal{out}}(\mathcal{A},P)}{\gamma_{n^{\prime}+1}},$
and note that we must have $a_{n}\rightarrow 0$ since we have assumed $\beta_{1,n-1}^{\textnormal{out}}(\mathcal{A},P)/\gamma_{n}\rightarrow 0$. Then applying Lemma 5.2, for any $m$ we have
$\frac{\beta_{m,n-1}^{\textnormal{out}}(\mathcal{A},P)}{\gamma_{n}}\leq\frac{\sum_{k=n-1}^{n+m-2}\beta_{1,k}^{\textnormal{out}}(\mathcal{A},P)}{\gamma_{n}}\leq\sum_{k=n-1}^{n+m-2}\frac{\beta_{1,k}^{\textnormal{out}}(\mathcal{A},P)}{\gamma_{k+1}}\leq ma_{n},$
where the first step applies the lemma (and the assumption $\gamma_{n}>0$), the second step holds since we have assumed $\gamma_{n}$ is a nonincreasing sequence, and the last step holds by definition of $a_{n}$. Now, since $a_{n}\to 0$, we can fix a sequence $m_{n}\to\infty$ with $m_{n}\leq\left\lceil a_{n}^{-1/2}\right\rceil$. Then we have
$\frac{\beta_{m_{n},n-1}^{\textnormal{out}}(\mathcal{A},P)}{\gamma_{n}}\leq m_{n}a_{n}\leq a_{n}^{1/2}+a_{n}.$
[…]
Finally, defining
[…]
we see that $\epsilon_{n},\delta_{n}\rightarrow 0$.
This completes the proof of the first part of Theorem 3.6, i.e., asymptotic coverage for the inflated jackknife+ interval. The proof of the second claim in Theorem 3.6, i.e., for the uninflated jackknife+ interval, is similar; we give the details in the Appendix. For full conformal, the proof of Theorem 3.7 (proving coverage of the inflated version of the full conformal prediction set), follows via an identical argument as for the inflated jackknife+, i.e., by combining the finite-sample result of Theorem 3.5 with Lemma 5.2; we omit the details. ∎
@@ 5.3 An empirical comparison
As mentioned above, the upper bound given in Lemma 5.2, which effectively tells us that $1$-stability at level $\beta$ implies $m$-stability at level $m\beta$, is loose—it is often the case that the actual out-of-sample or in-sample $m$-stability of an algorithm is far better than this coarse upper bound. To see this, we run a simulation to examine the out-of-sample stability for four algorithms: $k$-nearest neighbors, ridge regression, decision trees, and random forest. For $k$-nearest neighbors, the upper bound is actually fairly tight: we see only a mild gap between the actual value of $\beta_{m,n}^{\textnormal{out}}$ and the upper bound provided by the lemma. For the other three algorithms, however, the gap is substantial, showing that the upper bound of the lemma can be quite pessimistic (and in particular, showing that the algorithm $\mathcal{A}$ may indeed lead to strong training-conditional coverage results, according to our theorems).
@@ 5.4 Examples of $m$-stable algorithms: theoretical results
In this section, we will examine three types of regression algorithms, to establish theoretical bounds on their $m$-stability. In particular, we will see that, for some of them, the scaling as $m$ increases behaves more favorably than the coarse bound given above in Lemma 5.2. This parallels some of the findings in our simulation in Section 5.3 above, which showed several practical algorithms that exhibit good $m$-stability, beyond the coarse bound predicted by Lemma 5.2. All proofs for results in this section are deferred to the Appendix.
@@ 5.4.1 $k$-nearest neighbors
In particular, this result suggests that (for $m\leq n$) we have $\beta_{m,n}(\mathcal{A},P)=\mathcal{O}(\frac{m}{n})$ for both in- and out-of-sample stability, which means that $\beta_{m,n}(\mathcal{A},P)\propto m\cdot\beta_{1,n}(\mathcal{A},P)$, as suggested by the upper bound proved in Lemma 5.2. This agrees with the findings of our simulation in Section 5.3, where we observed that the bound given in Lemma 5.2 is relatively tight for the $k$-nearest neighbors algorithm (unlike the other algorithms).
@@ 5.4.3 Bagging
The following proposition extends this result to $m$-stability for a general $m$. (For simplicity, we only consider out-of-sample stability in this result; similar techniques can also establish slightly weaker bounds for $\beta_{m,n}^{\textnormal{in}}$.)
@@
If we take $\frac{N}{n-N}=\mathcal{O}(1)$ as is typically the case in practice (e.g., by choosing $N=n/2$), we therefore see that $\beta_{m,n}^{\textnormal{out}}(\mathcal{A}_{\textnormal{bag}},P)=\mathcal{O}(\sqrt{\frac{m}{n}}+\frac{1}{\sqrt{B}})$, meaning that we may have $\beta_{m,n}^{\textnormal{out}}(\mathcal{A}_{\textnormal{bag}},P)\ll m\cdot\beta_{m,n}^{\textnormal{out}}(\mathcal{A}_{\textnormal{bag}},P)$, particularly if $B$ is small—in other words, the upper bound of Lemma 5.2 may be loose. This is supported by our simulation results in Section 5.3 for the random forest algorithm (which is a specific instance of bagging—namely, by taking the base algorithm $\mathcal{A}$ to be a regression tree—although in that case, resampled data sets are constructed by sampling with replacement rather than without replacement).
@@ 6 Proofs of main results
6.1.1 Step 1: training-conditional coverage versus coverage on the test set.
First we will relate the training-conditional coverage $\alpha_{P}^{\textnormal{$\gamma$-J+}}(\mathcal{D}_{n})$ to coverage for the test set. Let
@@
6.1.2 Step 2: define an oracle.
@@
6.1.3 Step 3: relating the jackknife+ to the oracle.
@@
$\mathbf{1}\left\{\textnormal{Quantile}\left(\{\gamma+R_{j}-R_{i;j}\}_{j\in[n]};1-\alpha\right)<0\right\}=1,\quad\mathbf{1}\left\{\textnormal{Quantile}\left(\{\widetilde{R}_{j}-\widetilde{R}_{i}\}_{j\in[n]};1-\alpha^{\prime\prime}\right)<0\right\}=0.$
@@
$\alpha_{P}^{\textnormal{J+}}(\mathcal{D}_{n})=\mathbb{P}\{Y\not\in\widehat{C}^{\textnormal{J+}}_{n}(X)\mid\mathcal{D}_{n}\}\\ =\mathbb{P}\{Y\not\in\widehat{C}^{\textnormal{$\gamma$-J+}}_{n}(X)\mid\mathcal{D}_{n}\}+\mathbb{P}\{Y\in\widehat{C}^{\textnormal{$\gamma$-J+}}_{n}(X)\backslash\widehat{C}^{\textnormal{J+}}_{n}(X)\mid\mathcal{D}_{n}\}\\ =\alpha_{P}^{\textnormal{$\gamma$-J+}}(\mathcal{D}_{n})+\mathbb{P}\{Y\in\widehat{C}^{\textnormal{$\gamma$-J+}}_{n}(X)\backslash\widehat{C}^{\textnormal{J+}}_{n}(X)\mid\mathcal{D}_{n}\},$
@@
6.3.1 Step 1: training-conditional coverage versus coverage on the test set.
@@
the $\gamma$-inflated full conformal prediction interval trained on $\{(X_{j},Y_{j})\}_{j\in[n]}$ and evaluated at a test point $x$, and define
@@
6.3.2 Step 2: define an oracle.
@@
6.3.3 Step 3: relating the full conformal interval to the oracle.
@@ Acknowledgements
R.F.B. was was partially supported by the Office of Naval Research via grant N00014-20-1-2337, and by the National Science Foundation via grant DMS-2023109. The authors thank Michael Bian for helpful feedback on an earlier draft of this manuscript.
Appendix A Additional proofs
A.1 Completing the proof of Theorem 3.6
In Section 5.2, we verified the asymptotic training-conditional coverage result for the inflated jackknife+ interval, i.e., the first claim of Theorem 3.6. To complete the proof, we now need to verify the second claim in Theorem 3.6, i.e., for the uninflated jackknife+ interval; this will follow from a very similar argument as for the first claim.
Redefining
$a_{n}=\max_{n^{\prime}\geq n-1}\beta^{\textnormal{out}}_{1,n^{\prime}}(\mathcal{A},P),$
we have $a_{n}\to 0$ due to the assumption $\beta^{\textnormal{out}}_{1,n}(\mathcal{A},P)\to 0$. Applying Lemma 5.2, then, for any $m$ we have
$\beta^{\textnormal{out}}_{m,n-1}(\mathcal{A},P)\leq\sum_{k=n-1}^{n+m-2}\beta^{\textnormal{out}}_{1,k}(\mathcal{A},P)\leq ma_{n}.$
Taking $m_{n}=\left\lceil a_{n}^{-1/2}\right\rceil$ as before, we again have $m_{n}\to\infty$ and $\beta^{\textnormal{out}}_{m_{n},n-1}(\mathcal{A},P)\leq a_{n}^{1/2}+a_{n}$. Applying Theorem 3.3 with $\delta=e^{-\sqrt{\min\{n,m_{n}\}}}$, and plugging in our bound on $\beta^{\textnormal{out}}_{m_{n},n-1}(\mathcal{A},P)$, we then have
@@ Appendix A Additional proofs
which completes the proof of the desired claim by taking
@@
and observing that $\epsilon_{n},\delta_{n}\to 0$, as before.
A.2 Proof of Theorem 4.1 (jackknife+—alternative version)

### v3

@@ 4.3 In-sample, out-of-sample, or neither
However, there is a third option: the test point might be in the training set for one model, but not the other. In particular, Barber et al. (2021) consider this alternative type of stability, which (adjusting to the notation of our present paper) can be expressed by the quantity
@@
We write “in/out” in this definition to emphasize the fact that, in this case, the test point $X_{n+1}$ is in the training sample for one of the models ($\widehat{\mu}_{n+1}$) but out of the training sample for the other ($\widehat{\mu}_{n}$). We remark that Barber et al. (2021) refer to this as “in-sample stability” but we change the terminology here to allow for all the various types of stability under consideration.
@@ 5.1 Relating $1$-stability and $m$-stability
At a high level, it is intuitive that achieving $m$-stability, for some $m>1$, is more challenging than achieving $1$-stability. For swap-stability, this intuition is correct, as shown in the following proposition.
@@
Fix any algorithm $\mathcal{A}$ and distribution $P$. For any $n\geq m^{\prime}>m>0$,
@@
On the other hand, if we define stability by removing/adding data points rather than by swapping, this may not necessarily hold, due to degenerate counterexamples (e.g., we can take an algorithm $\mathcal{A}$ which, when run on a training sample of size $n$, returns the constant function $\widehat{\mu}(x)\equiv C\cdot\textnormal{mod}(n,2)$ for an arbitrary constant $C>0$; then $\beta_{1,n}^{\textnormal{out}}(\mathcal{A},P)=\beta_{1,n}^{\textnormal{in}}(\mathcal{A},P)=C$, but $\beta_{m,n}^{\textnormal{out}}(\mathcal{A},P)=\beta_{m,n}^{\textnormal{in}}(\mathcal{A},P)=0$ for any even $m$). Nonetheless, in practice, we generally expect $\beta_{m,n}^{\textnormal{out}}(\mathcal{A},P)$ and $\beta_{m,n}^{\textnormal{in}}(\mathcal{A},P)$ to increase monotonically in $m$.
@@
For any algorithm $\mathcal{A}$ and any distribution $P$, the following inequality holds for out-of-sample stability for any $n,m\geq 0$:
@@
In particular, this lemma immediately implies the results of Theorems 3.6 and 3.7. (We defer the detailed proof to the Appendix A.) However, as we will demonstrate below both empirically (in Section 5.2) and theoretically (in Section 5.3), this result is fundamentally somewhat weak: it does not capture the favorable stability properties that appear to hold in practice, and we will see that the $m$-stability of algorithms typically tends to be much better than this coarse upper bound.
@@ 5.2 An empirical comparison
As mentioned above, the upper bound given in Lemma 5.2, which effectively tells us that $1$-stability at level $\beta$ implies $m$-stability at level $m\beta$, is loose—it is often the case that the actual out-of-sample or in-sample $m$-stability of an algorithm is far better than this coarse upper bound. To see this, we run a simulation to examine the out-of-sample stability for four algorithms: $k$-nearest neighbors, ridge regression, decision trees, and random forest. Figure 1 shows the simulation results. For $k$-nearest neighbors, the upper bound is actually fairly tight: we see only a mild gap between the actual value of $\beta_{m,n}^{\textnormal{out}}$ and the upper bound provided by the lemma. For the other three algorithms, however, the gap is substantial, showing that the upper bound of the lemma can be quite pessimistic and in particular, showing that the algorithm $\mathcal{A}$ may indeed lead to strong training-conditional coverage results, according to our theorems.
@@ 5.3 Examples of $m$-stable algorithms: theoretical results
In this section, we will examine three types of regression algorithms, to establish theoretical bounds on their $m$-stability. In particular, we will see that, for some of them, the scaling as $m$ increases behaves more favorably than the coarse bound given above in Lemma 5.2. This parallels some of the findings in our simulation in Section 5.2 above, which showed several practical algorithms that exhibit good $m$-stability, beyond the coarse bound predicted by Lemma 5.2.
@@ 5.3.1 $k$-nearest neighbors
In particular, this result suggests that for $m\leq n$, we have $\beta_{m,n}(\mathcal{A},P)=\mathcal{O}(\frac{m}{n})$ for both in- and out-of-sample stability, which means that $\beta_{m,n}(\mathcal{A},P)\propto m\cdot\beta_{1,n}(\mathcal{A},P)$, as suggested by the upper bound proved in Lemma 5.2. This agrees with the findings of our simulation in Section 5.2, where we observed that the bound given in Lemma 5.2 is relatively tight for the $k$-nearest neighbors algorithm (unlike the other algorithms).
@@ 5.3.3 Bagging
The following proposition extends this result to $m$-stability for a general $m$. For simplicity, we only consider out-of-sample stability in this result; similar techniques can also establish slightly weaker bounds for $\beta_{m,n}^{\textnormal{in}}$.
@@
If we take $\frac{N}{n-N}=\mathcal{O}(1)$ as is typically the case in practice (e.g., by choosing $N=n/2$), we therefore see that $\beta_{m,n}^{\textnormal{out}}(\mathcal{A}_{\textnormal{bag}},P)=\mathcal{O}(\sqrt{\frac{m}{n}}+\frac{1}{\sqrt{B}})$, meaning that we may have $\beta_{m,n}^{\textnormal{out}}(\mathcal{A}_{\textnormal{bag}},P)\ll m\cdot\beta_{m,n}^{\textnormal{out}}(\mathcal{A}_{\textnormal{bag}},P)$, particularly if $B$ is small—in other words, the upper bound of Lemma 5.2 may be loose. This is supported by our simulation results in Section 5.2 for the random forest algorithm, which is a specific instance of bagging—namely, by taking the base algorithm $\mathcal{A}$ to be a regression tree—although in that case, resampled data sets are constructed by sampling with replacement rather than without replacement.
@@ 6 Proofs of main results
6.1.1 Step 1: training-conditional coverage versus coverage on the test set
First we will relate the training-conditional miscoverage $\alpha_{P}^{\textnormal{$\gamma$-J+}}(\mathcal{D}_{n})$ to miscoverage for the test set. Let
@@ 6.1.1 Step 1: training-conditional coverage versus coverage on the test set
6.1.2 Step 2: define an oracle
@@
6.1.3 Step 3: relating the jackknife+ to the oracle
@@
$\mathbf{1}\left\{\textnormal{Quantile}\left(\{\gamma+R_{j}-R_{i;j}\}_{j\in[n]};1-\alpha\right)<0\right\}=1,$
and
$\quad\mathbf{1}\left\{\textnormal{Quantile}\left(\{\widetilde{R}_{j}-\widetilde{R}_{i}\}_{j\in[n]};1-\alpha^{\prime\prime}\right)<0\right\}=0.$
@@ 6.1.4 Combining everything
$\alpha_{P}^{\textnormal{J+}}(\mathcal{D}_{n})=\mathbb{P}\{Y\not\in\widehat{C}^{\textnormal{J+}}_{n}(X)\mid\mathcal{D}_{n}\}=\mathbb{P}\{Y\not\in\widehat{C}^{\textnormal{$\gamma$-J+}}_{n}(X)\mid\mathcal{D}_{n}\}+\mathbb{P}\{Y\in\widehat{C}^{\textnormal{$\gamma$-J+}}_{n}(X)\backslash\widehat{C}^{\textnormal{J+}}_{n}(X)\mid\mathcal{D}_{n}\}\\ =\alpha_{P}^{\textnormal{$\gamma$-J+}}(\mathcal{D}_{n})+\mathbb{P}\{Y\in\widehat{C}^{\textnormal{$\gamma$-J+}}_{n}(X)\backslash\widehat{C}^{\textnormal{J+}}_{n}(X)\mid\mathcal{D}_{n}\},$
@@
6.3.1 Step 1: training-conditional coverage versus coverage on the test set
@@
the $\gamma$-inflated full conformal prediction set trained on $\{(X_{j},Y_{j})\}_{j\in[n]}$ and evaluated at a test point $x$, and define
@@
6.3.2 Step 2: define an oracle
@@
6.3.3 Step 3: relating the full conformal interval to the oracle
@@ Acknowledgements
The authors thank the Associate Editor and anonymous referees for their valuable comments. The authors also thank Michael Bian for helpful feedback on an earlier draft of this manuscript. R.F.B. was partially supported by the Office of Naval Research via grant N00014-20-1-2337, and by the National Science Foundation via grant DMS-2023109.
Appendix A Proof of asymptotic results
Proof of Theorems 3.6 and 3.7.
First, define $m_{n}\leq\left\lceil(b_{n}/\gamma_{n})^{-1/2}\right\rceil$, and note that $m_{n}\to\infty$ and $m_{n}b_{n}/\gamma_{n}\to 0$ since $b_{n}/\gamma_{n}\to 0$. Define also
@@ Appendix A Proof of asymptotic results
Observe that $\epsilon_{n},\delta_{n}\rightarrow 0$, since $m_{n}\to\infty$ and $m_{n}b_{n}/\gamma_{n}\to 0$.
Next, for any algorithm $\mathcal{A}$ and any distribution $P$, if $\beta^{*,\textnormal{out}}_{1,n-1}(\mathcal{A},P)\leq b_{n}$, then applying Lemma 5.2 with $m=m_{n}$, we have
$\beta_{m_{n},n-1}^{\textnormal{out}}(\mathcal{A},P)\leq\sum_{k=n-1}^{n+m_{n}-2}\beta_{1,k}^{\textnormal{out}}(\mathcal{A},P)\leq m_{n}b_{n}.$
$\mathbb{P}\left\{\alpha_{P}^{\textnormal{$\gamma_{n}$-J+}}(\mathcal{D}_{n})\geq\alpha+\epsilon_{n}\right\}\leq\delta_{n},$
for all such $\mathcal{A},P$. In other words, we have shown that
$\sup\limits_{\begin{subarray}{c}\mathcal{A},P\textnormal{ s.t. }\\ \beta_{1,n-1}^{*,\textnormal{out}}(\mathcal{A},P)\leq b_{n}\end{subarray}}\mathbb{P}\left\{\alpha_{P}^{\textnormal{$\gamma_{n}$-J+}}(\mathcal{D}_{n})\geq\alpha+\epsilon_{n}\right\}\leq\delta_{n}.$
This completes the proof of the first part of Theorem 3.6, i.e., asymptotic coverage for the inflated jackknife+ interval.
We now proceed to verify the second claim in Theorem 3.6, i.e., for the uninflated jackknife+ interval; this will follow from a very similar argument as for the first claim.
First, define $m_{n}=\lceil(b_{n}B_{n})^{-1/2}\rceil$, and note that $m_{n}\to\infty$ and $m_{n}b_{n}B_{n}\to 0$ since $b_{n}B_{n}\to 0$. Define also
@@
Observe that $\epsilon_{n},\delta_{n}\rightarrow 0$, since $m_{n}\to\infty$ and $m_{n}b_{n}B_{n}\to 0$.
Next, for any algorithm $\mathcal{A}$ and any distribution $P\in\mathcal{P}_{B_{n}}$, if $\beta^{*,\textnormal{out}}_{1,n-1}(\mathcal{A},P)\leq b_{n}$, then applying Lemma 5.2 we have $\beta_{m_{n},n-1}^{\textnormal{out}}(\mathcal{A},P)\leq m_{n}b_{n}$, as before. Therefore, by Theorem 3.3 with $\delta=e^{-\sqrt{\min\{n,m_{n}\}}}$,
$\mathbb{P}\left\{\alpha_{P}^{\textnormal{J+}}(\mathcal{D}_{n})\geq\alpha+\epsilon_{n}\right\}\leq\delta_{n}.$
Since this holds for all such $\mathcal{A},P$, in other words we have shown that
$\sup_{\begin{subarray}{c}\mathcal{A},P\in\mathcal{P}_{B_{n}}\textnormal{ s.t.}\\ \beta^{*,\textnormal{out}}_{1,n-1}(\mathcal{A},P)\leq b_{n}\end{subarray}}\mathbb{P}\left\{\alpha_{P}^{\textnormal{J+}}(\mathcal{D}_{n})\geq\alpha+\epsilon_{n}\right\}\leq\delta_{n},$
which completes the proof of Theorem 3.6.
For full conformal, the proof of Theorem 3.7 follows via an identical argument as that for the jackknife+ with inflation, i.e., by combining the finite-sample result of Theorem 3.5 with Lemma 5.2; we omit the details. ∎
Appendix B Proofs of training-conditional coverage under alternative stability frameworks
B.1 Proof of Theorem 4.1 (jackknife+—alternative version)